Consider the following reaction: CH4(g) + 2 H2S(g) ⇌ CS2(g) + 4 H2(g)A reaction mixture initially contains 0.50 M CH4 and 0.75 M H2S. If the equilibrium concentration of CS2 is 0.15 M, what is the equilibrium constant (Kc) for the reaction.Hint: You might need an I.C.E. Table for this one.

please show work
Answer #1

CH 4(g) + 2H2S(g) CS 2(g) + 4H2(g)

Initial 0.50. 0.75 0 0

Equilibrium. 0.50-x 0.76-2x x 4x

Since [CS2] = x = 0.15

Hence, [CH4] = 0.50-x = 0.50-0.15 = 0.35

[H2S] = 0.76-2x = 0.76-2×0.15 = 0.76-0.30 = 0.46

[H2] = 4x = 4×0.15 = 0.60

Kc = [CS2][H2]4/[CH4][H2S]2

= 0.15×(0.60)4/(0.46)2(0.35)

= 0.26